OpenStudy (anonymous):

what is the surface area of the cube below? Will give medal and fan!!(:

OpenStudy (anonymous):

@Candy13106

OpenStudy (anonymous):

@ETG311

OpenStudy (anonymous):

Surface area of a cube is given by: SA = 6s^2 (Where s = sidelength)

OpenStudy (anonymous):

is that 6 or 8 on the side?

OpenStudy (anonymous):

8 on the side.

OpenStudy (anonymous):

I got b? idk there all 8 (: @ETG311

OpenStudy (anonymous):

So do: SA = 6(8)^2 = 6*64 = 384 cm^2

OpenStudy (anonymous):

So B is right :D

OpenStudy (anonymous):

Yea its B

OpenStudy (anonymous):

thank u guys! :D

OpenStudy (anonymous):

Your welcome......